You borrow $11,000 from the bank at an interest rate of 6%, compounded annually. You are required to make 10 equal end of-year payments to pay off the loan.

a) What is the amount of these equal payments?

b) What is the amount of the 10 payments if the first payment is not made until 3 years after receipt of the money?
Answer #1

Rate = 6%
PV = 11000
Number of Periods = 10

a) Amount of Equal Payments = 11000/(1-(1+6%)-10)/6% = 1494.55

b) Value of Loan amount after 3 years = 11000*(1+6%)3 = 13,101.18
Amount of Equal Payments =13,101.18/(1-(1+6%)-10)/6% = 1780.03

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well